OpenStudy (imtiaz7):

A body has weight 20N. How much force is required to move it vertically upward with an acceleration of 2m/s^2?

OpenStudy (imtiaz7):

@MelissaHolmes

OpenStudy (atlas):

The force should overcome the given weight and produce an acceleration of 2m/s^2

OpenStudy (atlas):

F2-F1 = m*a

OpenStudy (atlas):

F2 = F1 + m*a F2 = 20 + m *2

OpenStudy (atlas):

Now we know only the weight NOT Mass - and given weight you can calculate mass if you know the value of 'g' of the place

OpenStudy (atlas):

g =9.8 m/s^2 for earth and Weight = m*g